OpenStudy (anonymous):

given f(x)=3x^3+1x+2 , the value of the derivative of the INVERSE of the functioin at f^(-1) at x=-2

OpenStudy (anonymous):

ok i will show you in few steps how to solve this problem

OpenStudy (anonymous):

whats next?

OpenStudy (anonymous):

sorry let me do this again ok

OpenStudy (anonymous):

kk yeah i got the inverse im stuck after it

OpenStudy (anonymous):

wait is this x^3

OpenStudy (anonymous):

ok there is a different method then

OpenStudy (anonymous):

i will be right back sorry i got a phone call

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

back, sorry for the delay

OpenStudy (anonymous):

do you know anything about cubic formula?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i will do the whole answer but it will take a while typing it

OpenStudy (anonymous):

well this is the best method to apply in order to solve this problem

OpenStudy (anonymous):

if you dont mind, thanks:)

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

well it is simple, as it is a bit similar to quadratic equation but a bit more messy

OpenStudy (anonymous):

let's start

OpenStudy (anonymous):

Step one: rearrange the equation \[y = 3x^3+x+2\]

OpenStudy (anonymous):

\[x = 3y^3+y+2\]

OpenStudy (anonymous):

\[ 3y^3+y+(2-x)=0\]

OpenStudy (anonymous):

Step two: applying the cubic root formula \[ax^3+bx^2+cx+d=0\]

OpenStudy (anonymous):

\[x = \sqrt[3]{\frac{- b^3 }{ 27a^3 }+\frac{ bc }{ 6a^2 }-\frac{ d }{2a }+\sqrt{(\frac{- b^3 }{ 27a^3 }+\frac{ bc }{ 6a^2 }-\frac{ d }{2a })^2+(\frac{ c }{ 3a }-\frac{ b^2 }{ 9a^2 })^3}}\]

OpenStudy (anonymous):

\[+ \sqrt[3]{\frac{- b^3 }{ 27a^3 }+\frac{ bc }{ 6a^2 }-\frac{ d }{2a }+\sqrt{(\frac{- b^3 }{ 27a^3 }+\frac{ bc }{ 6a^2 }-\frac{ d }{2a })^2+(\frac{ c }{ 3a }-\frac{ b^2 }{ 9a^2 })^3}}-\frac{ b }{ 3a }\]

OpenStudy (anonymous):

these two posts are one equation

OpenStudy (anonymous):

now what are the coefficients of our third order polynomial

OpenStudy (anonymous):

\[a=3, b=0, c=1, d=2\]

OpenStudy (anonymous):

now in the above equation all the b in it are canceled which will simplify our expression that yields to the following :

OpenStudy (anonymous):

\[y = \sqrt[3]{\frac{ -d }{ 2a }+\sqrt{(\frac{ -d }{ 2a })^2+(\frac{ c }{ 3a })^3}}\]

OpenStudy (anonymous):

\[+\sqrt[3]{\frac{ -d }{ 2a }+\sqrt{(\frac{ -d }{ 2a })^2+(\frac{ c }{ 3a })^3}}\]

OpenStudy (anonymous):

again the last two posts are one equation separated by the plus sign

OpenStudy (anonymous):

recall \[a=3, b=0, c=1, d=(2-x)\]

OpenStudy (anonymous):

plug in those values

OpenStudy (anonymous):

do you know how to substitute those values in the above equation

OpenStudy (anonymous):

i will leave the rest to you, i have to fix my customers computer, then i will be back as soon as i can

OpenStudy (anonymous):

okay thanks